Daily Activities

From LCLS-II testing
Jump to: navigation, search

2/7/2019 - LERF Buncher & Quarter turned on! [1]

1/24/2019 - Software complete for Coriolis flow meters [2]

1/18/2019 - LCLS-II cryomodules are at 2K!!! [3]

1/17/2019 - MO is reactivated [4]

1/15/2019 - Permission granted to begin cooldown of LCLS-II CMs [5]

1/14/2019 - Check list complete [6], OSP signed off [7]

1/9/2019 - SLAC folks are here to commission magnet racks [8]

12/19/2018 - Demag coils in place & diode thermometers being wired [9]

12/18/2018 - Helium piping being plumbed in bridging area [10]

12/17/2018 - Cryo group has concerns about a single valve between clean & 'dirty' helium [11]

12/12/2018 - Beam tube connection being made between the J-12 & J-5 [12], also work progresses on transfer line welding [13]

12/11/2018 - Coriolis flow meters final wiring connections [14] [15]

12/10/2018 - Preparation for beam line connection in clean room [16]

11/21/2018 - Pressure test passed, on to final leak checks [17]

11/20/2018 - Coriolis flow meters are mounted & ready for cables, Cables are being prepped my fab folks [18]

11/19/2018 - Vault was cleaned up & remaining parts sorted for final push

11/18/2018 - The two cables for the Cryo-Can thermometry have been terminated on both ends, as soon as welding is complete these will be tested. [19]

11/16/2018 - The clean room is finished [20]

11/14/2018 - All LLRF cables are complete [21] [22] [23] [24]

11/9/2018 - Cryo-can is prepped for grouting [25], Diodes are shorted out with wire & aluminum foil prior to welding [26], Welding began on tying the cryo-can to the transferline [27]

11/8/2018 - Cryomodule J05 was moved closer to the end can so electricians an assemble mid point clean room [28]

11/8/2018 - Dimensional checks revealed cryo-can's side port was off by ~ 1cm, & end flange as well but the bellows is a bit long - so the dimensional stack-up is correct [29]

11/7/2018 - Bellows for connecting CM2 to the cryo-can was moved to provide access to the Coriolis flow meters [30] [31]

11/6/2018 - Cryo-can surveyed into it's final location

11/5/2018 - Cryo-can was moved to the vault

10/18/2018 - We met to discuss positioning on the cryomodule from position 1 to position 2, follow on meeting will be Monday to go over draft procedures

10/16/2018 - Cryo-can side bellows was delivered to LERF vault

10/2/2018 - Hotwire anemometer parts are nearly finished

8/16/2018 - Eden Cryo stated in the status call this AM stated that the acceptance testing will take place after Labor Day, the 10 flux gate magnetometers have arrived

8/15/2018 - 8 Gauge 2 conductor arrived for SC magnet connections [32]

8/14/2018 - The load test of the trolley beam has been delayed until the chain fall arrives in a couple weeks

8/13/2018 - Trolley beam has been mounted to the ceiling, trolley/chain-fall is on order [33]

8/11/2018 - Installation crew ready to drill the trolley beam anchor holes [34]

8/10/2018 - Rail is painted (way off white 8-) and ready for installation [35]

8/9/2018 - Trolley rail welding and inspection complete [36]

8/8/2018 - Trolley beam is ready for welding [37]

8/6/2018 - Degausing power supply from FermiLab has arrived and is unpacked - [38]

8/3/2018 - Welding, inspection, and leak checking is complete for the 2k/40k supply lines

8/1/2018 - Welding has begun on the field fit joint of the transferline [39]

7/31/2018 - Task list has been updated with trolley installation added [40]

7/25/2018 - Transferline is being positioned [41] [42]. Leak in CM2 water header has been repaired [43]

7/22/2018 - Nearly all of the cables are terminated in the vault for CM # 1

7/19/2018 - Clearance shown for cryo-can; looks like removing vent stack leaves 80.14" [44] Side view of Cryo-Can showing ~2" spacing for grouting in base plate [45]

Vault clearances, Overhead door - 95 1/2", Copper Vent line - 91", Waveguide for CM #2(easy to remove one section) - 83 1/2", New wire basket cable tray (easily raised) - 83"

7/18/2018 - Transferline have been delivered to the LERF vault [46] [47]

7/13/2018 - Travel to Eden Cryogenics for final inspection of transferline, [48], leak check has passed [49], pressure test has also passed [50], progress is also being made on the cryo-can [51] [52] [53]

7/10/2018 - Four sets of wheels have been delivered to enable shuffling of the cryomodules [54]

6/29/2018 - Isolators (3 pcs) have been inspected per request by Dian/SLAC, no anomalies were found in the three inspected.

6/21/2018 - Progress is made on the transfer lines [55] [56] [57]

6/20/2018 - Support structure for de-gaussing coils have been delivered to the vault

6/11/2018 - All 4 LLRF racks are now powered up

6/5/2018 - Coriolis flow meter wiring modifications finished [58] [59]

6/4/2018 - The last of the anchors have been placed for the raised floor, RF cable terminations continue for SSA drive & LO

6/2/2018 - Leak checking of the waveguide is in progress

6/1/2018 - Cables have begun to be dressed into the tray in the vault

5/31/2018 - The last pieces of cable tray have been installed in the vault

5/24/2018 - A master spreadsheet has been made to track all of the mechanical parts as they arrive

5/22/2018 - RF cable installation has begun

5/21/2018 - Waveguide in zone 1 (CM #1) has been leak checked successfully [60]

5/20/2018 - Raised floor sections have been cut, clearing the way for RF cable installation

5/19/2018 - JLab Open House!

5/14/2018 - Crew is getting JT actuators mounted for continuing cryo controls check out

5/12/2018 - Crew is in to continue cable pulls & Kevin & Dayne are landing cables and beginning commissioning activities on the cryo controls, waveguide is now complete upstairs with the final SSA connections [61]

5/11/2018 - Kevin & Dayne fully trained up - Thanks to David Hamlett for giving rad-con practical on Friday

5/10/2018 - Kevin & Dayne from SLAC have arrived to begin landing cables in the cryo control racks

5/7/2018 - Most of the ~80 cryo cables have been installed [62], these need to be dressed in the tunnel

5/5/2018 - Happy Cinco de Mayo! Cable pulls have begun 8-) [63] [64]

5/4/2018 - LLRF and Cryo-racks are moved into place [65] [66] [67] [68] [69] [70]

5/3/2018 - Dark light solenoid is moved to storage [71]

5/2/2018 - Cryo racks & LLRF were uncrated and craned into the LERF [72] [73] [74]

5/1/2018 - Crane arrived but w/o truck The 16th SSA and the two magnet racks were craned into LERF and then crane drove off... [75], SLAC shipment arrived

4/30/2018 - SLAC shipment did not arrive!

4/26/2018 - vault clean-up has been done, clearing path to remove MIT dark light solenoid from the vault

4/25/2018 - Water connections are being done on the SSAs and isolators. FL02 girders have been connected both up and down stream

4/24/2018 - Adding a final piece of cable tray in the area of the waveguide drops to the input couplers to enable proper dressing of the CM control cables, design work on cool down line continues [76] More images of the waveguide [77], [78]

4/23/2018 - Waveguide taps for PSS pressure interlock are being machined, once complete all connections in the gallery will be complete.

4/19/2018 - Cable label listing is ready [79] First 8 isolators are installed [80] Waveguide is complete for first SSA upstairs - the significance is that the design is now validated for all upstairs waveguide [81]

4/18/2018 - Lake LERF! Facilities have been contacted [82] Clean room for connecting beam line vacuum has been delivered to receiving

4/17/2018 - SLAC shipment arrives!

4/16/2018 - Stands are surveyed in and ready to grout - the first one will not be grouted to enable the installation if new CMs - that stand is in thre way of moving the CM around the corner [83]

4/13/2018 - Second set of stands (CM #1) have been installed, survey crew is positioning them [84]

4/12/2018 - First LCLS-II cryomodule has been moved to the LERF vault [85] [86] [87]

4/11/2018 - F100 cryomodule has been relocated to the north access service building for this summer's installation [88]

4/5/2018 - Base plates for CM #2 installed [89]

4/4/2018 - Base plates delivered [90]

4/2/2018 - Draft of cable pull listing released [91]

3/29/2018 - PSS AC power for SSAs is being checked out with electricians

3/26/2018 - Water manifolds for SSAs & circulators mounted [92]

3/6&7/2018 - Facilities Advisory Committee meeting at SLAC. Took advantage of the meeting to meet with folks from all of the systems that we will be getting operational at the LERF; High Power RF, Low Level RF, Cryogenic controls, Vacuum Controls, and Magnet controls as well as their associated software contact.

3/3/2018 - Waveguide installation party! Thanks Guys!

3/2/2018 - Control power is turned on the CM #1 SSAs, IP addresses are being set.

3/1/2018 - 3 Coriolis mass flow meters have been delivered, Installation instructions are found here; [93]

2/28/2018 - Waveguide work began on upstairs components

2/27/2018 - Network connections have been established for many of the devices in the gallery, here is link to IP addresses: [94]

2/26/2018 - Balance of waveguide components have been shipped - all ~400 pieces will be on site this week, now the isolators are pacing the gallery work. Field wiring terminal blocks finished for Cryo-Cons, "fake" RTDs have been installed in the blocks and controllers have been setup for 100 ohm RTDs. Feb. 26 , 2018 P&ID release

2/23/2018 - Waveguide connections for CM #2 are finished in the tunnel up to the directional couplers

2/22/2018 - Cryo-Con RTD controllers have been mounted in the rack FL03B24

2/21/2018 - Survey group has laid out grid from mounting base plates

2/15/2018 - Center connection heat shields have arrrived

2/9/2018 - Most all of remaining cable has been ordered

2/8/2018 - Electricians are working on the PSS panel, AC power for SSAs in zone 4

2/7/2018 - Crew is beginning to bolt waveguide together

2/6/2018 - PSS chassis wiring nearly complete for zone 3 (CM #1) [95], AC power run to SSAs in zone 3 [96]

2/5/2018 - Magnet water lines have been moved out of the way of waveguide installation

2/4/2018 - Fly Eagles Fly! Congrats on Super Bowl victory!

2/2/2018 - Solid State Amplifiers are being positioned in the gallery [97], Waveguide installation has begun in the tunnel [98]

2/1/2018 - Radiation assessment complete [99], SSAs are craned into gallery, 19 pin & 8 pin connectors have been delivered

1/31/2018 - Happy Super Blue Blood Moon! Installation crew has completed waveguide support structure

1/30/2018 - Servers installed in the gallery racks, 100% of waveguide strait sections have been delivered, an update on the project status was presented to the joiont call [100]

1/29/2018 - First racks are powered up in gallery

1/26/2018 - CryoCon temperature controllers have been delivered [101], server rack has power, 450 Waveguide gaskets were delivered [102], all 4 pin connectors were delivered [103]

1/25/2018 - Installation crew began drilling holes, mounting anchors, and unistrut for waveguide installation. [104], I hosted a tour for the SLAC folks, Excess cable was inventoried in transport trailers

1/24/2018 - PDFs of waveguide layout is posted [105] [106] [107], Survey is finished with waveguide support layout in the vault, clean room ordered for making center connection between CMs [108]

1/23/2018 - I hosted Elvin Harms for a visit, survey crew is marking vault ceiling for the waveguide supports, racks were jacked up

1/22/2018 - Waveguide deliveries have been inventoried and separated out to the gallery & vault locations, 8 more SSAs were delivered

1/19/2018 - MS connectors for CMs have been ordered, AC power work continues

1/18/2018 - More waveguide delivered, rack ready for jacking, threaded rod support for waveguide layout being surveyed

1/17/2018 - New AC panel mounted next to E1000, this will supply 120 VAC for all new racks and equipment.

1/16/2018 - Cryomodule in zone 2 is being surveyed into place.

1/15/2018 - Happy MLK Day!

1/12/2018 - Waveguide supports have been installed in both zones [109], PSS box mounted

1/11/2018 - Ac Power requirements have been posted [110], the rack prep is complete - they are ready to be jacked up, Dark Light power supply moved to west side door for extraction.

1/10/2018 - Klystron racks are ready to be jacked up

1/4/2018 - Happy Snow Day 8 !

12/22/2017 - C37 moved back into place completing the shuffle! (placed 'order' with C. Curtis for final alignment...) [111]

12/21/2017 - Began placing orders for Cryogenic controls and Vacuum system cables

12/19/2017 - Rack bases were delivered

12/18/2017 - Zone 2 base plates have been grouted this AM [112], CryoCon 18i cryodiode controllers have been ordered [113], PC/PLC installed and communicating [114]

12/15/2018 - Clean room draft design sent to manufacturer for quote [115], Dell servers have been delivered!

12/14/2017 - Third review was held for the cryo-Can assembly that ties the CMs to the cryo transfer lines Mounting base plates have been relocated, and surveyed in for C37, zone 2. These are ready to re-grout. PCs are ready for installation & network connection - these are remote accessed machines with no local keyboard & monitor

12/13/2017 - PCs arrived for PLC remote programming, PLC arrived from SLAC to be programed

12/12/2017 - Nuts & bolts for waveguide installation have been delivered

12/6/2017 - Wheels are attached to zone 3 for movement to south wall for storage

12/5/2017 - Girder 1F04 removed for ease of CM move, waveguide gaskets ordered

12/4/2017 - Zone 4 klystrons removed

11/28/2017 - JLab electricians have been trained to remove Klystrons

11/22/2017 - 7/8" Heliax for Total Loss Monitor delivered

11/21/2017 - Waveguide order awarded

11/20/2017 - Waveguide order tweaked to reflect delivery times for SLAC type mitered H & E bends vs standard parts

11/17/2017 - Kevin visited Mega Ind., the waveguide company

11/15/2017 - 7/8" Heliax ordered for Total Loss Monitor (TLM)

11/14/2017 - Heliax PR submitted - 6,000' 3/8"

11/13/2017 - Second Cryo Can design review [116]

11/10/2017 - Waveguide bids have been received

11/9/2017 - Network switch installed and operational in Zone 3 FL03B13

11/8/2017 - Kevin visiting FermiLab - discussions with OPs, CM Assembly, Interlock group & others

11/7/2017 - Semi-annual Personnel Safety System re-certification complete

11/6/2017 - Draft P & ID drawing done & posted [117]

11/3/2017 - GW magnets were radcon surveyed and moved to storage in BLDG 36, 2 PCs were ordered for PLC programming of Cryo controls & vacuum controls, service request submitted to move rack from free stock to LERF gallery - this is a 36" deep rack for the servers (& free {;-), the network has been created for the project; lclsnet - [118], Active VVUs have been installed on 480 VAC panels [119]

11/2/2017 - All of the girders have been removed in preparation of CM move & shuffle [[120]]

11/1/2017 - Last of the cryomodule vacuums & girders have been secured & the clean room has been removed from the vault, the cryomodule & waveguide layout has been taped out on vault floor [[121]]

10/31/2017 - Cable length & routing checks being made

10/30/2017 - Network switches ordered, 2 racks moved to gallery, cryomodule vacuum isolation continues

10/27/2017 - 3 of the 6 vacuum connections for the CEBAF style cryomodules have been secured, had a cable planning meeting

10/26/2017 - The downstream side of zone 2 has gone through the 'slow' bleed-up process

10/25/2017 - Vacuum has been secured from the upstream side of Zone 2 CM in the vault

10/24/2017 - 480 VAC panels are being mounted in gallery

10/11/2017 - Waveguide design review

10/10/2017 - Clean Hoods have been moved to vault

10/9/2017 - Electricians have "Air Gapped" all power connections to FL03 & FL04 CPS & HPAs